# A Stitch in Time: Learning Procedural Workflow via Self-Supervised Plackett–Luce Ranking

### v1

@@ D.1 Five-Fold Cross-Validation on Surgical Phase Recognition
In the surgical domain, robust evaluation is critical to ensure that models generalize effectively across varying patient anatomies and surgical workflows. To verify the statistical robustness of our method in this challenging setting, we performed 5-fold cross-validation across all three surgical datasets: Cholec80 [Twinanda2017EndoNet:Videos], AutoLaparo [Wang2022AutoLaparo:Hysterectomy], and M2CAI16 [Stauder2016TheChallenge]. We report the linear probing top-1 accuracy (Mean $\pm$ Std) to demonstrate the stability of the learned features.
@@ D.2 Sensitivity Analysis of t-SNE Visualization
In the qualitative evaluation presented in the main manuscript, we employed a default t-SNE perplexity of 30, which offers a balanced representation of both local and global feature structures. To verify that the observed class separability is an intrinsic property of the learned embeddings rather than an artifact of visualization parameter tuning, we provide a robustness analysis on the Cholec80 dataset [Twinanda2017EndoNet:Videos] in Fig. 9.
@@ D.3 Additional Attention Maps
We provide an extended qualitative analysis of the attention focus of the model by visualizing self-attention maps queried by the [CLS] token across diverse procedural contexts in Fig. 10. This comparison encompasses both surgical scenes from the AutoLaparo dataset [Wang2022AutoLaparo:Hysterectomy] and cooking activities from the GTEA dataset [Fathi2011LearningActivities]. PL-Stitch consistently concentrates high attention weights within task-relevant areas and demonstrates a strong semantic alignment with the workflow. For instance, in the AutoLaparo examples, our PL-Stitch model’s attention remains anchored on the instrument-tissue interaction sites and demonstrates robust tracking of the surgical flow. Similarly, in the GTEA examples, attention accurately tracks the manipulated objects and active interaction zones, such as the condiment container and the spread on the bread. In contrast, baseline methods such as DINO, iBOT, and VideoMAEv2 exhibit significantly more diffuse attention patterns that often drift towards background elements or fail to distinctly highlight the active interaction site. This comparison underscores the stability and precision of PL-Stitch in localizing key visual cues compared to prior self-supervised approaches.

### v2

@@ D.1 Five-Fold Cross-Validation on Surgical Phase Recognition
In the surgical domain, robust evaluation is critical to ensure that models generalize effectively across varying patient anatomies and surgical workflows. To verify the statistical robustness of our method in this challenging setting, we performed 5-fold cross-validation across all three surgical datasets: Cholec80 [47], AutoLaparo [54], and M2CAI16 [41]. We report the linear probing top-1 accuracy (Mean $\pm$ Std) to demonstrate the stability of the learned features.
@@ D.2 Sensitivity Analysis of t-SNE Visualization
In the qualitative evaluation presented in the main manuscript, we employed a default t-SNE perplexity of 30, which offers a balanced representation of both local and global feature structures. To verify that the observed class separability is an intrinsic property of the learned embeddings rather than an artifact of visualization parameter tuning, we provide a robustness analysis on the Cholec80 dataset [47] in Fig. 9.
@@ D.3 Additional Attention Maps
We provide an extended qualitative analysis of the attention focus of the model by visualizing self-attention maps queried by the [CLS] token across diverse procedural contexts in Fig. 10. This comparison encompasses both surgical scenes from the AutoLaparo dataset [54] and cooking activities from the GTEA dataset [12]. PL-Stitch consistently concentrates high attention weights within task-relevant areas and demonstrates a strong semantic alignment with the workflow. For instance, in the AutoLaparo examples, our PL-Stitch model’s attention remains anchored on the instrument-tissue interaction sites and demonstrates robust tracking of the surgical flow. Similarly, in the GTEA examples, attention accurately tracks the manipulated objects and active interaction zones, such as the condiment container and the spread on the bread. In contrast, baseline methods such as DINO, iBOT, and VideoMAEv2 exhibit significantly more diffuse attention patterns that often drift towards background elements or fail to distinctly highlight the active interaction site. This comparison underscores the stability and precision of PL-Stitch in localizing key visual cues compared to prior self-supervised approaches.